OpenStudy (anonymous):

Solve the equation. y + (–6) = –9 A. –15 B. –3 C. 54 D. 3

OpenStudy (anonymous):

y=-9-6 y=-3

OpenStudy (anonymous):

is it b -3

OpenStudy (uri):

-9-6=-15..not -3 @fluffy95